# Joint Constraints on Fuzzy and Warm Dark Matter from Satellite Populations of the Milky Way and Andromeda

### v1

@@ 5.2 Warm dark matter
Figure 4 shows the marginalized PDFs and contour maps of the shared parameters in the WDM model. The marginalized posterior yields $M_{\mathrm{hm}}<3.20\times 10^{7}~M_{\odot}$ at 95% CL, corresponding to a lower limit on the WDM particle mass using Equation (11), and we have
$m_{\mathrm{WDM}}>6.22~\mathrm{keV}\ (95\%\ \mathrm{CL}).$ (20)
@@
This constraint represents a $\sim$5% improvement over the result obtained using only the MW satellite population by 2025ApJ...986..127N with $m_{\mathrm{WDM}}>5.9~\mathrm{keV}$. At a posterior ratio of 20:1, we obtain $M_{\mathrm{hm}}<4.24\times 10^{7}~M_{\odot}$, corresponding to
@@
Since the posterior distributions for all parameters in the WDM case are very similar to those in the FDM case, we do not provide figures showing the posterior distributions of all the MW and M31 parameters for the WDM case. We also rerun our model using the larger M31 mass prior from 2023ApJ...948..104P, obtaining $m_{\mathrm{WDM}}>6.10~\mathrm{keV}$ (95% CL) and $m_{\mathrm{WDM}}>5.63~\mathrm{keV}$ (20:1 posterior ratio).
@@
We note that in 2025ApJ...986..127N, the inferred $M_{\mathrm{hm}}$ for WDM is slightly lower than that for FDM, whereas our results show the opposite trend. This difference also explains why our improvement in the WDM constraint is only marginal. A possible reason is that our number of realizations, $N_{\mathrm{real}}$, is much larger than that used in 2025ApJ...986..127N, making our model more sensitive to subtle differences in the subhalo population. As shown in Figure 1, at the most sensitive mass scale around $10^{8}\,M_{\odot}$, the subhalo mass function for WDM is slightly higher than that for FDM at the same $M_{\mathrm{hm}}$. To compensate for this difference, the inferred $M_{\mathrm{hm}}$ for WDM tends to be slightly larger, resulting in the extended tail as shown in Figure 4.
@@
Compared to 2025ApJ...986..127N, our improvements arise mainly from two aspects: first, we adopt a more flexible mass-scaling approach; second, we include the M31 satellite population in our joint analysis. Since the constraints are primarily sensitive to the faintest satellites, incorporating the M31 satellite population strengthens the limits only at the few-percent level, which we summarize as an overall improvement of $\sim 10\%$ relative to using the MW sample alone. Moreover, because the mass and accretion history of M31 are highly uncertain, we do not treat all parameters as shared parameters, as was done in 2024ApJ...967...61N, when presenting our final constraints. We have verified that treating all parameters as shared would further tighten the constraints by about 5% relative to our adopted approach.
@@
Figure 5: Luminosity functions of satellites located between 30 and 300 kpc from the centers of the MW (blue) and M31 (gray), based on the FDM results from our modeling framework. Dark and light shaded regions show the 68% and 95% CLs, and the black solid lines denotes the mean predicted luminosity functions.
6 Additional Results and Discussion
6.1 Luminosity functions of the MW and M31
Using the posterior distributions inferred from our modeling framework, and adopting the FDM results as a representative case, we predict the satellite luminosity functions for both the MW and M31 (see Figure 5). Within the range $20<M_{V}<-5.5$ and distances of $30~\mathrm{kpc}<R<300~\mathrm{kpc}$ from the centers of the MW or M31, we obtain $42\pm 7$ satellites for the MW and $110\pm 33$ for M31. These results are consistent with the estimates from 2025arXiv250912313T, who found $29^{+9}_{-7}$ MW satellites, and from 2023ApJ...952...72D, who obtained $136^{+65}_{-35}$ or $92^{+19}_{-26}$ M31 satellites, depending on the assumed spatial distribution.
@@ 6.1 Luminosity functions of the MW and M31
We find that the number of M31 satellites is approximately three times that of the MW, which is consistent with the findings given by 2025arXiv250912313T. Considering that satellite abundance is known to scale roughly linearly with host halo mass, this may suggest that M31 is about three times more massive than the MW, in agreement with the estimate from 2023ApJ...948..104P. However, the enhanced satellite abundance around M31 may also be partly explained by its past interaction with M33 (2013MNRAS.430...37C; 2025arXiv250206948D).
6.2 Anisotropy in the M31 satellite distribution
@@
As more M31 satellites have been discovered, it has become evident that their spatial distribution may be not isotropic. Specifically, among the 24 M31 satellites used in this work, only three are located on the side opposite the MW (2022ApJ...938..101S; 2023ApJ...952...72D). To examine whether this anisotropy could arise from observational selection effects, we use the posterior distributions inferred from our modeling framework to simulate $10^{5}$ realizations, each representing an observation of 24 M31 satellites.
@@
We then obtain the PDF and cumulative distribution function (CDF) of the number of satellites located opposite the MW (see Figure 6). In all $10^{5}$ simulated realizations, none produce as few as three satellites on the far side. Even the probability of having eight satellites on the far side corresponds to a CDF value below $0.05$. This indicates that, even after accounting for distance uncertainties, the observed distribution cannot be explained by selection effects alone, consistent with the conclusions of previous studies (2022ApJ...938..101S; 2023ApJ...952...72D; 2025NatAs...9..692K). Interestingly, although detectability decreases with increasing distance, the larger observable volume counteracts this decline. As a result, if the M31 satellite distribution is isotropic, the probability of detecting more satellites on the far side than on the near side would in fact be slightly greater than 50%. Finally, we note that the strong anisotropy of the M31 satellites may, in principle, affect abundance-based constraints, but no reliable framework currently exists to incorporate this effect. In our analysis, we do not include any spatial information in our binning scheme. We have also verified that using different simulated host halos yields consistent constraints within our modeling framework. We will explore this issue in more detail in future work.
@@
7 Summary and Outlook
In this work, we develop a modeling framework and perform a joint analysis of the MW and M31 satellite systems to study the properties of both FDM and WDM. We use the galaxy–halo connection model together with the updated transfer functions and subhalo mass functions to construct the framework, and adopt the observational data of MW and M31 satellite galaxies from DES, PS1, and PAndAS, including their corresponding selection functions. We further account for the uncertain virial masses of the MW and M31 by introducing priors on their host halo masses and linearly scaling the relevant model parameters.
We find that, for the FDM case, $m_{\mathrm{FDM}}>1.75\times 10^{-20}~\mathrm{eV}$ (95% CL) and $m_{\mathrm{FDM}}>1.41\times 10^{-20}~\mathrm{eV}$ (20:1 posterior ratio). For thermal-relic WDM, we find $m_{\mathrm{WDM}}>6.22~\mathrm{keV}$ (95% CL) and $m_{\mathrm{WDM}}>5.75~\mathrm{keV}$ (20:1 posterior ratio). These results represent a moderate improvement over the MW-only constraints given by the previous works. Our analysis therefore provides the strongest constraints to date on FDM and WDM models derived from satellite galaxy populations in the Local Group. We also verify that adopting extreme assumptions for the mass of M31 weakens our constraints only slightly. These limits can be combined with complementary probes such as strong gravitational lensing and the Lyman-$\alpha$ forest to achieve even tighter bounds (2021ApJ...917....7N; 2021MNRAS.506.5848E).
Looking ahead, next-generation surveys such as the Rubin Observatory Legacy Survey of Space and Time (LSST; 2025OJAp....8E..89T), Chinese Space Station Survey Telescope (CSST; 2023MNRAS.523..876Q; 2025arXiv250704618C), and Dark Energy Spectroscopic Instrument (DESI; 2023ApJ...944....1D) will greatly expand the known satellite populations in the Local Group, including systems fainter than those currently detected. In addition, these surveys will also improve measurements of the virial masses and accretion histories of the MW and M31. Combined with advances in high-precision, large-volume dark-matter simulations that probe a broader range of host masses and assembly histories, as well as continued progress in observational selection functions, future analyses will be able to achieve more accurate constraints on the properties of dark matter.
We thank Ethan Nadler for the help with implementing the galaxy–halo connection model code. J.X.L. and Y.G. acknowledge the support from National Key R&D Program of China grant Nos. 2022YFF0503404 and 2020SKA0110402, and the CAS Project for Young Scientists in Basic Research (No. YSBR-092). KL was supported by National Key R&D Programme of China (no. 2024YFC2207400). This work is also supported by science research grants from the China Manned Space Project with grant Nos. CMS-CSST-2025-A02, CMS-CSST-2021-B01, and CMS-CSST-2021-A01.
@@ Appendix A Posterior distributions of model parameters
In Figures 7 and 8, we show the PDFs and contour maps (68% and 95% CLs) of the MW and M31 parameters for the FDM case, using the model introduced in Section 4. We do not show the $M_{\mathrm{host}}$ parameter because its posterior closely follows the adopted prior.
As discussed in Section 4.3, the posterior distributions of the M31-specific parameters inferred from our model are likely to be biased. This is because the virial mass of M31 is probably larger than the mass range of the simulated host halos we use, and its accretion history remains highly uncertain (2018ApJ...864...34S; 2019ApJ...872...24V; 2023ApJ...948..104P; 2024MNRAS.528.2653Z; 2025arXiv250206948D). As shown in Figures 7 and 8, the MW and M31 exhibit significant differences in three parameters: the power-law slope of the luminosity function, $\alpha$, the luminosity scatter, $\sigma_{M}$, and the amplitude of the galaxy–halo size relation, $\mathcal{A}$. This behavior is consistent with our expectations.
However, the faintest satellite in our M31 sample has an absolute magnitude of $M_{V}\simeq-6.0~\mathrm{mag}$, while the faintest MW satellite reaches $M_{V}\simeq 0~\mathrm{mag}$. The absolute magnitude of $M_{V}\simeq-6.0~\mathrm{mag}$ lies at the luminosity scale that is most sensitive to the reionization. Therefore, the fainter systems, which in our model appear only among the MW satellites, tend to form most of their stars before reionization. As a result, they develop more compact stellar distributions and deviate from the general size-luminosity relation (2022MNRAS.516.3944M; 2025arXiv250912313T; 2025arXiv250206948D). Their luminosity function may also deviate from a simple power law (2018ApJ...863..123B; 2022MNRAS.516.3944M). This effect could partially explain the differences in the inferred parameters between the MW and M31. To test this, we rerun our model using only MW satellites with $M_{V}<-6.0$. Although the parameter uncertainties increase due to the reduced data size, the discrepancy between the MW and M31 parameters decreases to some extent (with noticeable shifts in their maximum-posterior values), especially for the amplitude of the galaxy–halo size relation, $\mathcal{A}$.
Figure 7: The PDF and contour maps (68% and 95% CLs) of the MW parameters for the FDM case. The shaded regions in the one-dimensional marginalized PDFs and the parameter summaries denote 68% CL. The parameters $\sigma_{M}$, $\sigma_{\mathrm{gal}}$, and $\sigma_{\log_{10}R}$ are in dex, $\mathcal{M}_{50}$ and $M_{\mathrm{hm}}$ are in units of $M_{\odot}$, $\mathcal{A}$ is reported in pc, and $\alpha$, $n$, and $\mathcal{B}$ are dimensionless.

### v2

@@ V.2 Warm dark matter
Figure 4 shows the marginalized PDFs and contour maps of the shared parameters in the WDM model. The marginalized posterior yields $M_{\mathrm{hm}}<3.23\times 10^{7}~M_{\odot}$ at 95% CL, corresponding to a lower limit on the WDM particle mass using Equation (11), and we have
$m_{\mathrm{WDM}}>6.20~\mathrm{keV}\ (95\%\ \mathrm{CL}).$ (20)
@@
This constraint represents a $\sim$5% improvement over the result obtained using only the MW satellite population by E. O. Nadler et al. (2025) with $m_{\mathrm{WDM}}>5.9~\mathrm{keV}$. At a posterior ratio of 20:1, we obtain $M_{\mathrm{hm}}<4.23\times 10^{7}~M_{\odot}$, corresponding to
@@
Since the posterior distributions for all parameters in the WDM case are very similar to those in the FDM case, we do not provide figures showing the posterior distributions of all the MW and M31 parameters for the WDM case. We also rerun our model using the larger M31 mass prior from E. Patel & K. S. Mandel (2023), obtaining $m_{\mathrm{WDM}}>6.07~\mathrm{keV}$ (95% CL) and $m_{\mathrm{WDM}}>5.61~\mathrm{keV}$ (20:1 posterior ratio).
@@
We note that in E. O. Nadler et al. (2025), the inferred $M_{\mathrm{hm}}$ for WDM is slightly lower than that for FDM, whereas our results show the opposite trend. This difference also explains why our improvement in the WDM constraint is only marginal. A possible reason is that our number of realizations, $N_{\mathrm{real}}$, is much larger than that used in E. O. Nadler et al. (2025), making our model more sensitive to subtle differences in the subhalo population. As shown in Figure 1, at the most sensitive mass scale around $10^{8}\,M_{\odot}$, the subhalo mass function for WDM is slightly higher than that for FDM at the same $M_{\mathrm{hm}}$. To compensate for this difference, the inferred $M_{\mathrm{hm}}$ for WDM tends to be slightly larger, resulting in the extended tail as shown in Figure 4.
@@
Compared to E. O. Nadler et al. (2025), our improvements arise mainly from two aspects: first, we adopt a more flexible mass-scaling approach; second, we include the M31 satellite population in our joint analysis. Since the constraints are primarily sensitive to the faintest satellites, incorporating the M31 satellite population strengthens the limits only at the few-percent level, which we summarize as an overall improvement of $\sim 10\%$ relative to using the MW sample alone. This also explains why the results are insensitive to the methodological differences in selection functions and the inclusion or exclusion of the $M_{V}\simeq-6~\mathrm{mag}$ M31 satellite, since the M31 satellite population does not dominate the final constraints.
Moreover, because the mass and accretion history of M31 are highly uncertain, we do not treat all parameters as shared parameters, as was done in E. O. Nadler et al. (2024), when presenting our final constraints. We have verified that treating all parameters as shared would further tighten the constraints by about 5% relative to our adopted approach. However, as discussed above, enforcing the sharing of all parameters could introduce unphysical biases given the significant uncertainties in M31’s mass and accretion history. Therefore, this $\sim 5\%$ improvement likely stems from the compensation for these biased parameters rather than a genuine physical signal.
@@
Figure 5: Luminosity functions of satellites located within $300~\mathrm{kpc}$ from the centers of the MW (left panel) and M31 (right panel), derived from the FDM results in our modeling framework. The gray shaded regions represent the predicted luminosity functions, while the blue shaded regions show the predicted observable luminosity functions after applying the observational selection functions. The black stepped lines indicate the observed satellite counts from DES and PS1 (left) and PAndAS (right). Dark and light shaded regions correspond to the 68% and 95% CLs, respectively.
VI Additional Results and Discussion
VI.1 Luminosity functions of the MW and M31
Using the posterior distributions inferred from our modeling framework, and adopting the FDM results as a representative case, we predict the satellite luminosity functions for both the MW and M31. As shown in Figure 5, our model reproduces the observed luminosity functions of both the MW and M31 remarkably well, with the observational data falling within the 68% CLs. This validates the reliability of our modeling framework and the resulting constraints. The fact that the M31 luminosity function is well-reproduced further suggests that the significant uncertainties in M31’s halo mass and accretion history do not substantially bias our results. In other words, given the current observational precision, our constraints are relatively insensitive to the specific details of M31’s host halo properties. Furthermore, when performing the same check for a model where all parameters are shared between the MW and M31, we find that the MW luminosity function is slightly overestimated, while that of M31 is significantly underestimated. This discrepancy reinforces the conclusion that enforcing a fully shared parameter space yields unphysical results.
@@ VI.1 Luminosity functions of the MW and M31
For satellites with $M_{V}<0~\mathrm{mag}$, we predict a total of $222^{+47}_{-46}$ satellites within $300~\mathrm{kpc}$ of the MW, which is consistent with the result of E. O. Nadler et al. (2020). For satellites with $M_{V}<-5.5~\mathrm{mag}$, we predict $41^{+7}_{-7}$ and $80^{+24}_{-23}$ satellites within $300~\mathrm{kpc}$ of the MW and M31, respectively. These values are in good agreement with the results from C. Y. Tan et al. (2025b) and A. Doliva-Dolinsky et al. (2023). We find that the abundance of satellites in M31 is approximately twice that of the MW, a trend that was also noted by C. Y. Tan et al. (2025b). Considering that satellite abundance is known to scale roughly linearly with host halo mass, this may suggest that the halo mass of M31 is approximately twice that of the MW. However, the enhanced satellite abundance around M31 may also be partly explained by its past interaction with M33 (S. C. Chapman et al., 2013; A. Doliva-Dolinsky et al., 2025).
VI.2 Anisotropy in the M31 satellite distribution
@@
As more M31 satellites have been discovered, it has become evident that their spatial distribution may be not isotropic. Specifically, among the 24 M31 satellites used in this work, only three are located on the side opposite the MW (A. Savino et al., 2022; A. Doliva-Dolinsky et al., 2023). To examine whether this anisotropy could arise from observational selection effects, we use the posterior distributions inferred from our modeling framework to simulate $10^{5}$ realizations, each representing an observation of 24 M31 satellites.
@@
We then obtain the PDF and cumulative distribution function (CDF) of the number of satellites located opposite the MW (see Figure 6). In all $10^{5}$ simulated realizations, none produce as few as three satellites on the far side. Even the probability of having eight satellites on the far side corresponds to a CDF value below $0.05$. This indicates that, even after accounting for distance uncertainties, the observed distribution cannot be explained by selection effects alone, consistent with the conclusions of previous studies (A. Savino et al., 2022; A. Doliva-Dolinsky et al., 2023; K. J. Kanehisa et al., 2025). Interestingly, although detectability decreases with increasing distance, the larger observable volume counteracts this decline. As a result, if the M31 satellite distribution is isotropic, the probability of detecting more satellites on the far side than on the near side would in fact be slightly greater than 50%. Finally, we note that the strong anisotropy of the M31 satellites may, in principle, affect abundance-based constraints, but no reliable framework currently exists to incorporate this effect. In our analysis, we do not include any spatial information in our binning scheme. We have also verified that using different simulated host halos yields consistent constraints within our modeling framework. The fact that our model is in good agreement with the observed data (as shown in Figure 5) further reinforces the reliability of our current approach. We will explore this issue in more detail in future work.
@@
VII Summary and Outlook
In this work, we develop a modeling framework and perform a joint analysis of the MW and M31 satellite systems to study the properties of both FDM and WDM. We use the empirical galaxy–halo connection model described in E. O. Nadler et al. (2020) together with the updated transfer functions and subhalo mass functions to construct the framework, and adopt the observational data of MW and M31 satellite galaxies from DES, PS1, and PAndAS, including their corresponding selection functions. We further account for the uncertain virial masses of the MW and M31 by introducing priors on their host halo masses and linearly scaling the relevant model parameters.
We find that, for the FDM case, $m_{\mathrm{FDM}}>1.74\times 10^{-20}~\mathrm{eV}$ (95% CL) and $m_{\mathrm{FDM}}>1.42\times 10^{-20}~\mathrm{eV}$ (20:1 posterior ratio). For thermal-relic WDM, we find $m_{\mathrm{WDM}}>6.20~\mathrm{keV}$ (95% CL) and $m_{\mathrm{WDM}}>5.75~\mathrm{keV}$ (20:1 posterior ratio). These results represent a moderate improvement over the MW-only constraints given by the previous works. Our analysis therefore provides the strongest constraints to date on FDM and WDM models derived from satellite galaxy populations in the Local Group. We also verify that adopting extreme assumptions for the mass of M31 weakens our constraints only slightly. These limits can be combined with complementary probes such as strong gravitational lensing and the Lyman-$\alpha$ forest to achieve even tighter bounds (E. O. Nadler et al., 2021a; W. Enzi et al., 2021).
Looking ahead, next-generation surveys such as the Rubin Observatory Legacy Survey of Space and Time (LSST; K. Tsiane et al. 2025), Chinese Space Station Survey Telescope (CSST; H. Qu et al. 2023; CSST Collaboration et al. 2025), and Dark Energy Spectroscopic Instrument (DESI; A. Dey et al. 2023) will greatly expand the known satellite populations in the Local Group, including systems fainter than those currently detected. In addition, these surveys will also improve measurements of the virial masses and accretion histories of the MW and M31. Combined with advances in high-precision, large-volume dark-matter simulations that probe a broader range of host masses and assembly histories, as well as continued progress in observational selection functions, future analyses will be able to achieve more accurate constraints on the properties of dark matter.
We thank Ethan Nadler for the help with implementing the empirical galaxy–halo connection model code and for his helpful suggestions. J.X.L. and Y.G. acknowledge the support from the CAS Project for Young Scientists in Basic Research (No. YSBR-092), and National Key R&D Program of China grant Nos. 2022YFF0503404 and 2020SKA0110402. KL was supported by National Key R&D Programme of China (no. 2024YFC2207400). This work is also supported by science research grants from the China Manned Space Project with grant Nos. CMS-CSST-2025-A02, CMS-CSST-2021-B01, and CMS-CSST-2021-A01.
@@ Appendix A Posterior distributions of model parameters
In Figures 7 and 8, we show the PDFs and contour maps (68% and 95% CLs) of the MW and M31 parameters for the FDM case, using the model introduced in Section IV. We do not show the $M_{\mathrm{host}}$ parameter because its posterior closely follows the adopted prior.
As discussed in Section IV.3, the posterior distributions of the M31-specific parameters inferred from our model are likely to be biased. This is because the virial mass of M31 is probably larger than the mass range of the simulated host halos we use, and its accretion history remains highly uncertain (M. Semczuk et al., 2018; R. P. van der Marel et al., 2019; E. Patel & K. S. Mandel, 2023; X. Zhang et al., 2024; A. Doliva-Dolinsky et al., 2025). As shown in Figures 7 and 8, the MW and M31 exhibit significant differences in three parameters: the power-law slope of the luminosity function, $\alpha$, the luminosity scatter, $\sigma_{M}$, and the amplitude of the galaxy–halo size relation, $\mathcal{A}$. This behavior is consistent with our expectations.
However, the faintest satellite in our M31 sample has an absolute magnitude of $M_{V}\simeq-6.0~\mathrm{mag}$, while the faintest MW satellite reaches $M_{V}\simeq 0~\mathrm{mag}$. The absolute magnitude of $M_{V}\simeq-6.0~\mathrm{mag}$ lies at the luminosity scale that is most sensitive to the reionization. Therefore, the fainter systems, which in our model appear only among the MW satellites, tend to form most of their stars before reionization. As a result, they develop more compact stellar distributions and deviate from the general size-luminosity relation (V. Manwadkar & A. V. Kravtsov, 2022; C. Y. Tan et al., 2025b; A. Doliva-Dolinsky et al., 2025). Their luminosity function may also deviate from a simple power law (S. Bose et al., 2018; V. Manwadkar & A. V. Kravtsov, 2022). This effect could partially explain the differences in the inferred parameters between the MW and M31. To test this, we rerun our model using only MW satellites with $M_{V}<-6.0~\mathrm{mag}$. Although the parameter uncertainties increase due to the reduced data size, the discrepancy between the MW and M31 parameters decreases to some extent (with noticeable shifts in their maximum-posterior values), especially for the amplitude of the galaxy–halo size relation, $\mathcal{A}$.
Figure 7: The PDF and contour maps (68% and 95% CLs) of the MW parameters for the FDM case. The shaded regions in the one-dimensional marginalized PDFs and the parameter summaries denote 68% CL. The parameters $\sigma_{M}$, $\mathcal{S}_{\mathrm{gal}}$, and $\sigma_{\log_{10}R}$ are in dex, $\mathcal{M}_{50}$ and $M_{\mathrm{hm}}$ are in units of $M_{\odot}$, $\mathcal{A}$ is reported in pc, and $\alpha$, $n$, and $\mathcal{B}$ are dimensionless.